# Bayesian WeakS-to-Strong from Text Classification to Generation

### v2

@@ Appendix D Ablation Study on Probability Estimation for Weak Sequences
Appendix E Weights of Weak Models
The selection for weights of weak models was explored in this section, where experiments on average, dynamic and fixed weights were conducted. For average, the weight $\lambda_{i}$ in Eqn. (2) and Eqn. (4) is set to the same value for each weak model. That is to say, the losses are calculated for each weak model respectively and then averaged as the final loss. For fixed weight, the weights are set to a fixed value based on the performance of different weak models. For dynamic weight, for a certain sample, the confidence in the utterance of a weak model is treated as a weight for that weak model, considering the importance of different weak models could vary on different samples. The results are shown in Table 7. Compared to the average, fixed weight improves the average PGR by about 8% because it assigns weights based on the performance of different weak models, rather than just averaging. Dynamic weight shows a noticeable decline in performance compared to fixed weight. It may be because our proposed per-token target probability already provided the student model with fine-grained information about the reliability of the label, making dynamic weight entirely redundant.
@@ Appendix E Weights of Weak Models
Appendix F Performance of Joint Decoding
F.1 The Impact of the Quality of Weak Models
@@
F.2 The Impact of Beam Size
@@
Appendix G Regularising Term of EDL
@@
Appendix H Prompt Used in Experiment
@@
Appendix I Licenses for Existing Assets
The licenses for each asset used is listed below:
Model/Dataset License
GPT2-Large Modified MIT License
OPT-1.3B MIT license
Pythia-1.4B Apache 2.0
BLOOM-1B1 RAIL License v1.0
TinyLlama_v1.1 Apache 2.0
Qwen-7B Tongyi Qianwen LICENSE AGREEMENT
Llama2-7B Custom commercial license
SciQ CC BY-NC 3.0 DEED
SLURP CC BY 4.0

### v4

@@ Appendix E Ablation Study on Probability Estimation for Weak Sequences
Appendix F Weights of Weak Models
The selection for weights of weak models was explored in this section, where experiments on average, dynamic and fixed weights were conducted. For average, the weight $\lambda_{i}$ in Eqn. (2) and Eqn. (4) is set to the same value for each weak model. That is to say, the losses are calculated for each weak model respectively and then averaged as the final loss. For fixed weight, the weights are set to a fixed value based on the performance of different weak models (refer to C for specific value). For dynamic weight, for a certain sample, the confidence in the utterance of a weak model is treated as a weight for that weak model, considering the importance of different weak models could vary on different samples. The results are shown in Table 9. Compared to the average, fixed weight improves the average PGR by about 8% because it assigns weights based on the performance of different weak models, rather than just averaging. Dynamic weight shows a noticeable decline in performance compared to fixed weight. It may be because our proposed per-token target probability already provided the student model with fine-grained information about the reliability of the label, making dynamic weight entirely redundant.
@@ Appendix F Weights of Weak Models
Appendix G Performance of Joint Decoding
G.1 The Impact of the Quality of Weak Models
@@
G.2 The Impact of Beam Size
@@
Appendix H Regularising Term of EDL
@@
Appendix I Prompt Used in Experiment
@@
Appendix J Licenses for Existing Assets
Models and datasets we used in this work were all downloaded from HuggingFace website, expect SLURP which is downloaded from https://github.com/pswietojanski/slurp. The licenses and paths for each asset used is listed below:
Model/Dataset License Huggingface Path
GPT2-Large Modified MIT License openai-community/gpt2-large
OPT-1.3B MIT license facebook/opt-1.3b
Pythia-1.4B Apache 2.0 EleutherAI/pythia-1.4b
BLOOM-1B1 RAIL License v1.0 bigscience/bloom-1b1
TinyLlama_v1.1 Apache 2.0 TinyLlama/TinyLlama_v1.1
Qwen-7B Tongyi Qianwen LICENSE AGREEMENT Qwen/Qwen-7B
Llama2-7B Custom commercial license meta-llama/Llama-2-7b-hf
SciQ CC BY-NC 3.0 DEED allenai/sciq
SLURP CC BY 4.0 N/A
CosmosQA CC BY 4.0 allenai/cosmos_qa